### v2

# Gated Integration of Low-Rank Adaptation for Continual Learning of Large Language Models

Abstract
Continual learning (CL), which requires the model to learn multiple tasks sequentially, is crucial for large language models (LLMs). Recently, low-rank adaptation (LoRA), one of the most representative parameter-efficient fine-tuning (PEFT) methods, has gained increasing attention in CL of LLMs. However, most existing CL methods based on LoRA typically expand a new LoRA branch to learn each new task and force the new and old LoRA branches to influence old tasks equally, potentially leading to forgetting. In this work, we propose a new method, called gated integration of low-rank adaptation (GainLoRA), for CL of LLMs. GainLoRA expands a new LoRA branch for each new task and introduces gating modules to integrate the new and old LoRA branches. Furthermore, GainLoRA leverages the new gating module to minimize the influence from the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance. Experimental results on CL benchmarks demonstrate that GainLoRA outperforms existing state-of-the-art methods. Code is available at https://github.com/liangyanshuo/gainlora.
1 Introduction
Continual learning (CL), which requires the model to learn multiple tasks sequentially, is crucial for large language models (LLMs) shi2024continual . Specifically, although existing LLMs have demonstrated strong performance for a wide range of tasks brown2020language ; chowdhery2023palm ; touvron2023llama ; touvron2023llama2 ; zhang2022opt with extensive pre-trained knowledge and further fine-tuning strategies, they may lose knowledge acquired from old tasks when learning multiple tasks sequentially. This phenomenon, known as catastrophic forgetting luo2023empirical ; parisi2019continual ; DBLP:journals/pami/WangZSZ24 ; wang2023orthogonal , highlights the need for developing effective CL methods for LLMs. Existing CL methods can be categorized into two main categories. The first category DBLP:conf/iclr/RazdaibiedinaMH23 assumes that task identities are available during inference, while the second category liang2024inflora ; zhao2024sapt tackles a more difficult and practical setting where task identities are unavailable during inference.
Recently, low-rank adaptation (LoRA) hu2021lora , one of the most representative parameter-efficient fine-tuning (PEFT) methods, has gained increasing attention in the CL of LLMs bohaoscalable ; wang2023orthogonal . Specifically, by reparameterizing pre-trained weights in a low-rank form, LoRA updates only a limited number of parameters to adapt LLMs to a downstream task, making the fine-tuning process much more efficient than updating all parameters of LLMs han2024parameter . This efficiency also benefits CL, making LoRA increasingly popular in CL of LLMs.
Most existing CL methods based on LoRA liang2024inflora ; zhao2024sapt typically expand a new LoRA branch for learning each new task while freezing all old LoRA branches. In this way, they avoid forgetting caused by directly updating the LoRA parameters of old tasks. However, to handle the practical CL scenario where task identities are unavailable at inference time, existing methods liang2024inflora ; smithcontinual ; wang2023orthogonal based on LoRA integrate new and old LoRA branches through a simple addition. Consequently, they force the new and old LoRA branches to influence old tasks equally, which means that the new LoRA branch may cause a relatively large change in the model’s output on old tasks. This leads to forgetting and degrades the model’s overall performance in CL.
In this work, we propose a new method, called gated integration of low-rank adaptation (GainLoRA), for CL of LLMs. The contributions of GainLoRA are listed as follows:
  • GainLoRA expands a new LoRA branch to learn each new task and introduces gating modules to integrate the new and old LoRA branches.
  • GainLoRA leverages the new gating module to minimize the influence from the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance.
  • Experimental results on CL benchmarks show that GainLoRA outperforms existing state-of-the-art CL methods.
2 Related Work and Preliminaries
2.1 Related Work
Parameter-Efficient Fine-Tuning Parameter-efficient fine-tuning (PEFT) methods tune a limited number of parameters to adapt a pre-trained model for downstream tasks, showing much higher efficiency than tuning all the parameters of the pre-trained model, especially for LLMs zhangadaptive . For example, Adapter houlsby2019parameter modifies the model architecture by introducing trainable modules into Transformer layers and tunes these modules for downstream tasks. Prompt-tuning DBLP:conf/emnlp/LesterAC21 and Prefix-tuning DBLP:conf/acl/LiL20 insert learnable tokens into the input and tune them for downstream tasks. Low-rank adaptation (LoRA) hu2021lora reparameterizes the original model parameters with low-rank matrices and tunes these matrices for downstream tasks. Although tuning significantly fewer parameters than full fine-tuning, PEFT can achieve comparable performance to full fine-tuning across a wide range of natural language processing (NLP) tasks fu2022adapterbias ; hu2021lora ; DBLP:conf/nips/MahabadiHR21 ; zaken2022bitfit .
Continual Learning There are three main types of CL methods, categorized as regularization-based methods, memory-based methods, and expansion-based methods. Regularization-based methods jung2020continual ; kirkpatrick2017overcoming ; li2017learning incorporate a regularization term to mitigate catastrophic forgetting. Memory-based methods DBLP:conf/nips/dAutumeRKY19 ; DBLP:conf/nips/LiangL23 ; lopez2017gradient ; sun2019lamol ; zhao2024sapt utilize memory mechanisms to preserve knowledge from old tasks. Expansion-based methods DBLP:conf/nips/Hung0WCCC19 ; li2019learn ; liang2023adaptive ; rusu2016progressive mitigate catastrophic forgetting by introducing new parameters for learning new tasks while typically freezing old parameters.
Many CL methods aranilearning ; li2017learning ; liang2023adaptive are designed to train models from scratch. Recent studies liang2024inflora ; DBLP:conf/iclr/RazdaibiedinaMH23 ; wang2025self ; wang2023orthogonal ; wang2022learning have shown that leveraging PEFT strategies to fine-tune pre-trained models enables CL methods to achieve superior performance across tasks. For example, some methods DBLP:conf/iclr/QinJ22 ; DBLP:conf/iclr/RazdaibiedinaMH23 ; wang2022learning ; zhao2024sapt use prompt-tuning for continual learning. They either maintain independent prompts for each task or maintain a prompt pool and perform query-key matching to learn new tasks. There are also many methods liang2024inflora ; smith2023construct ; wang2023orthogonal ; zhao2024sapt adopting LoRA for continual learning. Most of these methods expand a new LoRA branch to handle each new task while freezing old LoRA branches to mitigate catastrophic forgetting. However, they force the new and old LoRA branches to influence old tasks equally, potentially leading to forgetting.
2.2 Preliminaries
Problem Definition We follow existing CL works wang2023orthogonal ; zhao2024sapt to formalize the problem definition for CL of LLMs. Specifically, in CL, a sequence of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{T}\}$ is presented to the model sequentially, where $T$ denotes the total number of tasks. The $t$-th task $\mathcal{T}_{t}$ consists of a training dataset $\mathcal{D}_{t}$. For any given sample $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{D}_{t}$, $\bm{x}_{t}$ denotes an input sentence and $\bm{y}_{t}$ denotes the corresponding output. When learning the $t$-th new task, the model is required to mitigate catastrophic forgetting of the t1t-1 previously learned tasks.
Similar to existing CL works for LLMs bohaoscalable ; zhao2024sapt , we consider a more challenging CL setting with three key challenges: (1) the model is presented with a sequence of tasks spanning various types, such as dialogue generation, information extraction and so on; (2) the model is not provided with task identities at inference time; (3) the model must learn without access to real or synthetic samples from previously learned tasks.
Low-Rank Adaptation LoRA hu2021lora is a widely adopted PEFT method used for fine-tuning various pre-trained models, particularly LLMs. Specifically, let $\bm{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ represent a pre-trained weight in LLMs, where $d_{in}$ and $d_{out}$ are the input and output dimensions, respectively. LoRA introduces an additional branch consisting of two matrices, $\bm{A}\in\mathbb{R}^{d_{out}\times r}$ and $\bm{B}\in\mathbb{R}^{r\times d_{in}}$, where $r\ll{\rm min}(d_{in},d_{out})$. LoRA then modifies the forward propagation of this layer as
$\bm{e}=(\bm{W}+\bm{A}\bm{B})\bm{h}.$
Here, $\bm{h}$ and $\bm{e}$ denote the input and output, respectively. $\bm{A}$ is initialized to $\bm{0}$, and $\bm{B}$ is initialized with a Gaussian distribution. During fine-tuning for downstream tasks, the pre-trained weight $\bm{W}$ remains frozen, and only the parameters $\bm{A}$ and $\bm{B}$ are fine-tuned.
Figure 1: (a) shows the expandable LoRA architecture of our GainLoRA for learning the $t$-th new task. (b) shows that for each task $\mathcal{T}_{i}$, GainLoRA uses an independent gating module $g_{i}(\cdot)$ to generate integration coefficient $a_{i}$.
3 Methodology
Our GainLoRA employs an expandable LoRA architecture, which is illustrated in Figure 1 (a). Specifically, before learning the $t$-th task ($1\leq t\leq T$), GainLoRA first expands the LoRA architecture by introducing the $t$-th new branch with matrices $\bm{A}_{t}\in\mathbb{R}^{d_{out}\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times d_{in}}$. The new and old LoRA branches are then integrated as
$\bm{W}_{t}=\bm{W}_{t-1}+a_{t}\bm{A}_{t}\bm{B}_{t}=\sum_{i=1}^{t}a_{i}\bm{A}_{i}\bm{B}_{i},$ (1)
where $a_{i}$ is an integration coefficient that determines the influence of the $i$-th LoRA branch to the input $\bm{h}$. Note that $\bm{W}_{t-1}$ is a zero matrix when $t=1$. As a result, the forward propagation in this layer is modified as
$\bm{e}=(\bm{W}+\bm{W}_{t})\bm{h}.$ (2)
Finally, only the new LoRA branch (i.e. the $t$-th LoRA branch) is updated for the $t$-th new task, while all the old LoRA branches are frozen. After learning the $t$-th task, (2) is also used for inference across all test samples, thereby ensuring compatibility with the scenario where task identities are unavailable during inference.
Many existing CL methods based on LoRA liang2024inflora ; smith2023construct ; smithcontinual ; wang2023orthogonal ; zhao2024sapt share a similar architecture to our method, as illustrated in Figure 1 (a). However, these methods fix all coefficients $\{a_{i}\}_{i=1}^{t}$ in (1) to $1$, forcing the new and old LoRA branches influence old tasks equally. As a result, the new LoRA branch introduces a change of $\bm{A}_{t}\bm{B}_{t}\bm{h}$ to the output for inputs $\bm{h}$ associated with old tasks, potentially leading to forgetting. Although some methods attempt to mitigate this forgetting by imposing regularization smithcontinual or orthogonality constraints liang2024inflora on the new LoRA branch, the fixed integration coefficients $\{a_{i}\}_{i=1}^{t}$ still limit their performance, as demonstrated by the experimental results presented in Section 4. The method in zhao2024sapt does not force the new and old LoRA branches to influence old tasks equally but relies on replaying synthetic old samples to mitigate forgetting, making it unsuitable for the scenario considered in this work.
Different from existing methods, GainLoRA introduces an independent gating module $g_{i}(\cdot)$ for each task $\mathcal{T}_{i}$ to generate the integration coefficients ($1\leq i\leq T$). To mitigate the forgetting caused by the new task, GainLoRA leverages the gating module to minimize the influence from the new LoRA branch to the old tasks. The details will be introduced in the following subsections.
3.1 Architecture of Gating Modules
As illustrated in Figure 1 (b), given an input sample $\bm{x}$, the gating module $g_{i}(\cdot)$ generates the integration coefficient for the $i$-th LoRA branch, denoted as $a_{i}=g_{i}(\bm{x})$. The computation of $g_{i}(\cdot)$ is defined as
$\bm{p}_{i,0}$ $=\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})),$
$\bm{p}_{i,l}$ $=\sigma(\bm{G}_{i,l}\bm{p}_{i,l-1}),~l\in\{1,2,...,L\},$
$g_{i}(\bm{x})$ $=f(\bm{G}_{i,L+1}\bm{p}_{i,L}).$ (3)
Here, ${\rm Token}(\cdot)$ represents the tokenizer used in LLMs to extract token embeddings from the input $\bm{x}$. ${\rm Pool}(\cdot)$ denotes an average pooling operation applied to the token embeddings to produce a fixed-size vector. $\sigma(\cdot)$ denotes the non-linear activation function. $\bm{G}_{i,l}$ denotes the weight matrix for the $l$-th layer of $g_{i}(\cdot)$ ($1\leq l\leq L+1$). In the final layer, $\bm{G}_{i,L+1}$ is a vector that maps the input vector $\bm{p}_{i,L}$ to a scalar. Following existing works with gating mechanisms cho2014properties ; DBLP:journals/neco/HochreiterS97 , the function $f(\cdot)$ is designed to map a scalar to a value within $[0,1]$, that is, $f(\cdot):\mathbb{R}\rightarrow[0,1]$.
Note that the input to gating modules is the same as that of LLMs, denoted as $\bm{x}$, which differs from the input to LoRA in a specific layer, denoted as $\bm{h}$. During the learning of the $t$-th new task, only the new gating module $g_{t}(\cdot)$ is updated, while all the old gating modules $\{g_{i}(\cdot)\}_{i=1}^{t-1}$ remain frozen.
3.2 Minimizing the Influence from the New LoRA Branch to Old Tasks
GainLoRA minimizes the influence from the new LoRA branch to old tasks by making $a_{t}=g_{t}(\bm{x})$ as close to 0 as possible for any input $\bm{x}$ from old tasks $\{\mathcal{T}_{i}\}_{i=1}^{t-1}$. However, since we focus on the scenario where no real or synthetic samples from old tasks are accessible, directly optimizing $g_{t}(\bm{x})$ to 0 is impractical. To overcome this challenge, GainLoRA imposes constraints on the new gating module $g_{t}(\cdot)$, implicitly guiding $g_{t}(\bm{x})$ to close to 0 and reduce the influence of the new LoRA branch to old tasks.
In the following two subsections, we first describe the constraints imposed on the new gating module $g_{t}(\cdot)$ and explain how these constraints guide $g_{t}(\bm{x})$ close to 0 for any $\bm{x}$ from the old tasks. Then, we detail the implementation of these constraints during training.
3.2.1 Constraints on New Gating Module
To formalize the constraints imposed on the new gating module $g_{t}(\cdot)$, we define the subspace spanned by the inputs to $\bm{G}_{t,l}~(1\leq l\leq L+1)$ from the previous t1t-1 tasks as:
[…]
Note that subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ cannot be obtained directly due to the unavailability of samples from old tasks. However, by introducing additional constraints, $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ can be solved iteratively, which will be discussed in Section 3.2.2.
Initialization Constraints Before learning the $t$-th task, the following constraints are imposed on the initialization of the new gating module $g_{t}(\cdot)$:
${\rm Init}(\bm{G}_{t,L+1})$ $\bot\mathcal{M}_{t,L+1},~f(0)=0,$ (5)
where ${\rm Init}(\bm{G}_{t,L+1})$ denotes the initialization of $\bm{G}_{t,L+1}$. These constraints ensure that for any sample $\bm{x}$ from the old tasks, the integration coefficient satisfies
$a_{t}=g_{t}(\bm{x})=f({\rm Init}(\bm{G}_{t,L+1})\bm{p}_{t,L})=0,$ (6)
where $\bm{p}_{t,L}$ is defined in (3.1). The second equality holds since $\bm{G}_{t,L+1}={\rm Init}(\bm{G}_{t,L+1})$ before learning the $t$-th new task. The third equality holds because $f(0)=0$ and $\bm{p}_{t,L}\in\mathcal{M}_{t,L+1}$ for any $\bm{x}$ from previous t1t-1 tasks.
Updating Constraints During the learning of the $t$-th task, the following constraints are imposed on the updates to the new gating module $g_{t}(\cdot)$:
$\Delta\bm{G}_{t,l}\bot\mathcal{M}_{t,l}\quad\text{for}\quad 1\leq l\leq L+1,$ (7)
where $\Delta\bm{G}_{t,l}$ denotes the update to $\bm{G}_{t,l}$. Based on existing studies wang2021training ; liang2023adaptive , the constraints in (7) ensure that $g_{t}(\bm{x})$ remains unchanged for inputs $\bm{x}$ from the old tasks during the learning of the $t$-th task. Formally, the following proposition holds:
Proposition 3.1.
If the constraints in (7) are satisfied, subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task. Furthermore, for any input $\bm{x}$ from the previous t1t-1 tasks, $g_{t}(\bm{x})$ remains unchanged during the learning of the $t$-th task.
The proof of this proposition is provided in Appendix A.3. Since the initialization constraints in (5) ensure $g_{t}(\bm{x})=0$ before learning the $t$-th new task, $g_{t}(\bm{x})=0$ is preserved throughout the learning process if the updating constraints in (7) are satisfied.
The fact that subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged, as stated in Proposition 3.1, is essential for implementing the orthogonal constraints in (7). Specifically, as will be detailed in Section 3.2.2, orthonormal bases for the subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ are learned to enforce the orthogonal constraints in (5) and (7). Since the subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task, their orthonormal bases also remain unchanged, allowing them to be pre-computed before learning the $t$-th task, thus facilitating the implementation of orthogonal constraints in (5) and (7) throughout the learning process.
3.2.2 Implementation of Constraints
There exist many functions $f(\cdot):\mathbb{R}\rightarrow[0,1]$ satisfying $f(0)=0$. In this work, we define $f(\cdot)$ as
$f(b)=|2\cdot{\rm sigmoid}(b)-1|,$ (8)
where ${\rm sigmoid}(\cdot)$ denotes the sigmoid function. Other functions $f(\cdot):\mathbb{R}\rightarrow[0,1]$ that satisfy $f(0)=0$ are also applicable, and experiments with different choices of $f(\cdot)$ are provided in Appendix C.3.1. Better model performance can be expected by designing more effective $f(\cdot)$, but this is not the focus of this paper.
Implementing the orthogonal constraints in (5) and (7) is challenging due to the lack of samples from previous t1t-1 tasks to approximate the subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$. To address this issue, we further impose the following constraints on the initialization of $\bm{G}_{t,l}~(1\leq l\leq L)$:
${\rm Init}(\bm{G}_{t,l})$ $\leftarrow\bm{G}_{t-1,l}.$ (9)
This strategy initializes the first $L$ layers of $g_{t}(\cdot)$ using the corresponding layers from the previous gating module $g_{t-1}(\cdot)$. As a result, the first $L$ layers of $g_{t}(\cdot)$ can be viewed as being initialized and starting their training at the beginning of the first task, continuing until the $t$-th task. Simultaneously, the first $L$ layers in $g_{i}(\cdot)$ serve as checkpoints, preserving the state of $g_{t}(\cdot)$ after learning the $i$-th task ($1\leq i\leq t$). At this time, we can use existing method gradient projection memory (GPM) DBLP:conf/iclr/SahaG021 to iteratively learn a set of matrices $\{\bm{M}_{t,l}\}_{l=1}^{L+1}$, where the columns of $\bm{M}_{t,l}$ contribute to a set of orthonormal bases of subspace $\mathcal{M}_{t,l}$. Details of GPM are provided in Appendix A.1. Then, before learning the $t$-th task, the following operation can be performed on ${\rm Init}(\bm{G}_{t,L+1})$:
${\rm Init}(\bm{G}_{t,L+1})\leftarrow{\rm Init}(\bm{G}_{t,L+1})-\bm{M}_{t,L+1}\bm{M}_{t,L+1}^{T}{\rm Init}(\bm{G}_{t,L+1}).$ (10)
At this time, we have
$\bm{M}_{t,L+1}^{T}({\rm Init}(\bm{G}_{t,L+1})-\bm{M}_{t,L+1}\bm{M}_{t,L+1}^{T}{\rm Init}(\bm{G}_{t,L+1}))$
$=$ $\bm{M}_{t,L+1}^{T}(\bm{I}-\bm{M}_{t,L+1}\bm{M}_{t,L+1}^{T}){\rm Init}(\bm{G}_{t,L+1})$
$=$ $(\bm{I}-\bm{M}_{t,L+1}^{T}\bm{M}_{t,L+1})\bm{M}_{t,L+1}^{T}{\rm Init}(\bm{G}_{t,L+1}).$ (11)
Since the columns of $\bm{M}_{t,L+1}$ form an orthonormal basis, we have $\bm{M}_{t,L+1}^{T}\bm{M}_{t,L+1}=\bm{I}$, which means $\bm{I}-\bm{M}_{t,L+1}^{T}\bm{M}_{t,L+1}=\bm{O}$. Therefore, the equation in (3.2.2) is equal to zero matrix $\bm{O}$. Note that $\mathcal{M}_{t,L+1}$ is spanned by the columns of $\bm{M}_{t,L+1}$, ${\rm Init}(\bm{G}_{t,L+1})$ satisfies the constraints in (5) after the operation in (10).
Similarly, during the learning of the $t$-th task, the following operation can be performed on $\{\Delta\bm{G}_{t,l}\}_{l=1}^{L+1}$:
$\Delta\bm{G}_{t,l}$ $\leftarrow\Delta\bm{G}_{t,l}-\bm{M}_{t,l}\bm{M}_{t,l}^{T}\Delta\bm{G}_{t,l}.$ (12)
With the same proving process in (3.2.2), we can show that the update in (12) allows the update $\{\Delta\bm{G}_{t,l}\}_{l=1}^{L+1}$ to satisfy the constraints in (7).
3.3 Updating the New LoRA Branch
Our GainLoRA aims to effectively integrate new and old LoRA branches while mitigating forgetting caused by the new LoRA branch on old tasks. Since GainLoRA does not impose specific update strategies for the new LoRA branch, it is inherently compatible with various existing CL methods that adopt similar LoRA architecture as our method and can update the new LoRA branch liang2024inflora ; smithcontinual ; wang2023orthogonal . Since these existing methods fix all integration coefficients $\{a_{i}\}_{i=1}^{t}$ to 1, combining our method with these existing methods can enhance their performance, as demonstrated in Section 4.
Algorithm 1 GainLoRA for Continual Learning
Input: The data of different tasks $\{\mathcal{D}_{t}\}_{t=1}^{T}$.
Output: Learned LoRA parameters $\{(\bm{A}_{i},\bm{B}_{i})\}_{i=1}^{T}$ and gating modules $\{g_{i}(\cdot)\}_{i=1}^{T}$.
for $t$ in $1:T$ do
  Expand the $t$-th new LoRA branch with $\bm{A}_{t}$ and $\bm{B}_{t}$;
  Impose initialization constraints on the new gating module $g_{t}(\cdot)$ by (8), (9) and (10);
  Integrate new and old LoRA branches by (1);
  for $\mathcal{B}_{t}\subseteq\mathcal{D}_{t}$ do
   Compute the loss in (13) and perform;
   Perform backward propagation to compute the update of the new LoRA branch and the new gating module;
   Impose updating constraints on the update of the new gating module by (7);
  end for
end for
3.4 Whole Process of GainLoRA
Algorithm 1 outlines the whole process of our GainLoRA. Before learning the $t$-th new task $\mathcal{T}_{t}$, GainLoRA first expands the LoRA architecture by introducing the $t$-th new branch with matrices $\bm{A}_{t}$ and $\bm{B}_{t}$. Simultaneously, a new gating module $g_{t}(\cdot)$ is initialized through the operations specified in (8), (10) and (9) to ensure that the initialization constraints in (5) are satisfied. The new and old LoRA branches are then integrated using (1), and the forward propagation is modified as (2).
During the learning of the $t$-th task $\mathcal{T}_{t}$ with the corresponding dataset $\mathcal{D}_{t}$, our method follows existing methods wang2023orthogonal ; zhao2024sapt and computes the loss for the new task through
$\mathcal{L}_{t}=\frac{1}{|\mathcal{D}_{t}|}\sum_{(\bm{x}_{t},\bm{y}_{t})\in\mathcal{D}_{t}}\sum_{j=1}^{|\bm{y}_{t}|}{\rm log}\left[P(y_{t,j}|\bm{x}_{t},y_{t,1},...,y_{t,j-1})\right],$ (13)
where $\bm{y}_{t}=[y_{t,1},y_{t,2},...,y_{t,|\bm{y}_{t}|}]$. Each time, GainLoRA samples a mini-batch $\mathcal{B}_{t}$ to minimize the loss in (13) by updating the new LoRA branch and the new gating module $g_{t}(\cdot)$. During this process, the projections defined in (12) are applied to the parameters of $g_{t}(\cdot)$, ensuring that the update constraints in (7) are satisfied.
GainLoRA introduces a new gating module for each new task, which incurs additional parameters and computational cost when combined with other methods. Section 4 will demonstrate that the trainable parameters added by GainLoRA are limited, making the number of trainable parameters in GainLoRA comparable to other methods. Additionally, Appendix C.1 and C.2 will demonstrate that the computational cost introduced by GainLoRA is minimal compared to the original LLMs.
4 Experiments
4.1 Experimental Settings
Datasets Following existing CL methods DBLP:conf/iclr/RazdaibiedinaMH23 ; wang2023orthogonal ; zhao2024sapt , we evaluate different methods on SuperNI DBLP:conf/emnlp/WangMAKMNADASPK22 and Long Sequence DBLP:conf/iclr/RazdaibiedinaMH23 benchmarks. SuperNI benchmark includes various types of NLP tasks, including dialogue generation, information extraction, question answering, summarization, and sentiment analysis. Following the protocols of existing method zhao2024sapt , three tasks are selected from each type, resulting in 15 tasks. These tasks are arranged into two different task sequences with different orders, referred to as Order 1 and Order 2. Long Sequence benchmark consists of 15 diverse classification tasks, which are similarly arranged into two task sequences with different orders, referred to as Order 3 and Order 4. More details about the benchmarks and task sequences are provided in Appendix B.
Evaluation Metric We use ${\rm A}_{j,i}$ to denote the model’s performance on the $i$-th task once the model learns the $j$-th task. Specifically, ${\rm A}_{j,i}$ represents accuracy for classification tasks and Rouge-L lin2004rouge for other types of tasks. Following traditional CL works chaudhry2020continual ; chaudhry2019tiny , we employ average performance (AP) and forgetting (FT) to evaluate the model’s performance. The formulas for these two metrics are defined as
${\rm AP}=\frac{1}{T}\sum_{i=1}^{T}{\rm A}_{T,i},~~~{\rm FT}=\frac{1}{T-1}\sum_{i=1}^{T-1}({\rm max}_{l\in\{1,2,...,T-1\}}{\rm A}_{l,i}-{\rm A}_{T,i}),$ (14)
where $T$ denotes the total number of tasks in the task sequence. AP evaluates the model’s final performance, and FT quantifies the forgetting.
Baselines We compare our method with state-of-the-art CL methods, including LFPT5 DBLP:conf/iclr/QinJ22 , EPI wang2023rehearsal , MIGU du2024unlocking , EWC kirkpatrick2017overcoming , TASL feng2024tasl , KIFLoRA feng2024kif , IncLoRA wang2023orthogonal , C-LoRA smithcontinual , O-LoRA wang2023orthogonal , and InfLoRA liang2024inflora . Additionally, we introduce a simple baseline called SeqLoRA, which does not expand new LoRA branches but sequentially updates old LoRA parameters for new tasks and lacks mechanism to mitigate forgetting. Note that many CL methods based on pre-trained models in CV focus on classification tasks, relying either on carefully designed classifiers mcdonnell2023ranpac ; zhang2023slca ; DBLP:conf/nips/ZhaoZYD024 or the [CLS] token in ViT le2024mixture ; smith2023coda ; wang2023hierarchical ; wang2024hide ; wang2022s ; wang2022dualprompt ; wang2022learning . In contrast, we follow existing works in NLP wang2023orthogonal ; zhao2024sapt and adopt next-token prediction to handle both classification and generation tasks, where models raffel2020exploring ; touvron2023llama2 lack a [CLS] token. Consequently, these CV methods are incompatible with our setting and cannot be directly compared. For completeness, we adapt some of them to our setup and report results in Appendix C.7.
Implementation Details Following existing CL works ouyang2022training ; wang2023orthogonal ; wei2021finetuned , all methods are implemented with instruction tuning ouyang2022training and optimized using AdamW DBLP:conf/iclr/LoshchilovH19 . To ensure fair comparisons, for all the methods based on LoRA, we follow existing CL methods hu2021lora ; wang2023orthogonal ; zhao2024sapt by incorporating the LoRA architecture into the query and value components of the multi-head attention mechanism in each Transformer block. Similar to the existing CL methods for LLMs wang2023orthogonal ; zhao2024sapt , we use T5 raffel2020exploring , Llama-2 touvron2023llama2 and Llama-3 dubey2024llama as the base architectures. Each experiment is repeated three times with different seeds, and the average result is reported. More details, such as the learning rate, batch size, and architecture of the gating modules in GainLoRA, are provided in Appendix B.2 and Appendix B.3.
Table 1: Results on different task sequences with T5-large model. Results of methods with are copied from existing paper zhao2024sapt .
Method Order 1 Order 2 Order 3 Order 4
[…]
LFPT5 DBLP:conf/iclr/QinJ22 39.03 10.87 29.70 20.72 66.62 14.57 67.40 13.20
EPI wang2023rehearsal - - - - 75.19 0.77 75.10 2.44
MIGU+FT du2024unlocking - - - - 71.30 11.39 69.05 14.06
EWC kirkpatrick2017overcoming 15.32 26.78 18.19 30.28 43.24 23.66 46.25 32.90
TaSL feng2024tasl 27.51 18.53 28.05 17.39 71.37 6.20 73.11 6.52
KIFLoRA feng2024kif 28.33 16.44 30.31 16.27 72.19 3.10 73.72 4.75
SeqLoRA 7.30 47.60 7.03 47.97 49.46 27.60 33.81 45.53
IncLoRA wang2023orthogonal 12.33 41.93 16.65 36.56 61.19 13.63 62.46 15.92
C-LoRA smithcontinual 22.69 24.25 32.81 11.60 66.83 8.64 61.86 14.18
O-LoRA wang2023orthogonal 26.37 19.15 32.83 11.99 70.98 3.69 71.21 4.03
GainLoRA (O-LoRA) 47.84 2.26 46.84 2.91 73.37 3.02 76.01 2.49
InfLoRA liang2024inflora 39.78 7.64 39.57 8.93 75.15 4.19 75.79 3.47
GainLoRA (InfLoRA) 46.21 2.40 46.44 2.61 78.01 0.77 77.54 1.25
4.2 Experimental Results
Compare with Existing Methods We first follow existing works du2024unlocking ; zhao2024sapt and evaluate different CL methods using T5-Large. Since our method does not impose specific update strategies for the new LoRA branch, we adopt the same update strategies as the two state-of-the-art methods, O-LoRA wang2023orthogonal and InfLoRA liang2024inflora . Note that these two methods leverage LoRA architecture similar to our method but fix all integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1. Details of these two methods are provided in Appendix A.2. We use GainLoRA (O-LoRA) and GainLoRA (InfLoRA) to respectively denote our methods adopting O-LoRA and InfLoRA to update the new LoRA branch. GainLoRA is also compatible with other methods that leverage expandable LoRA architecture shown in Figure 1 (a), and we give some results in Appendix C.5.
The results are shown in Table 1. As we can see, our methods GainLoRA (O-LoRA) and GainLoRA (InfLoRA) outperform O-LoRA and InfLoRA in both AP and FT, respectively. This improvement demonstrates that fixing all coefficients $\{a_{i}\}_{i=1}^{T}$ to 1 leads to forgetting on old tasks, thereby limiting the performance of O-LoRA and InfLoRA. By effectively mitigating this forgetting, GainLoRA (O-LoRA) and GainLoRA (InfLoRA) achieve superior performance. Furthermore, our methods consistently achieve the best performance across all task sequences.
Figure 2 illustrates the variation in the average performance across all learned tasks for different methods throughout the CL process. As shown, GainLoRA consistently outperforms the performance of O-LoRA and InfLoRA throughout the whole training process.
Figure 2: The variation of performance across different CL methods during training on different task sequences.
Table 2: The overall results on different task sequences with T5-XL model.
Method Order 1 Order 2 Order 3 Order 4
[…]
O-LoRA wang2023orthogonal 36.50 11.42 40.64 6.37 73.77 2.70 76.19 3.56
GainLoRA (O-LoRA) 50.10 3.21 49.86 3.04 78.41 2.59 77.21 3.30
InfLoRA liang2024inflora 45.61 5.60 45.85 5.10 80.22 2.09 79.43 1.71
GainLoRA (InfLoRA) 50.06 1.86 50.26 2.64 81.22 0.58 80.30 0.75
Table 3: The overall results on different task sequences with Llama-2-7B, Llama-2-13B and Llama-3-8B.
Models Methods Order 1 Order 2
[…]
Llama-2-7B O-LoRA wang2023orthogonal 39.37 15.84 37.55 20.23
GainLoRA (O-LoRA) 51.10 4.96 51.14 5.57
InfLoRA liang2024inflora 42.93 11.23 39.94 15.00
GainLoRA (InfLoRA) 51.27 2.84 50.17 4.71
Llama-2-13B O-LoRA wang2023orthogonal 43.92 14.15 40.05 19.53
GainLoRA (O-LoRA) 52.47 4.78 51.68 5.86
InfLoRA liang2024inflora 43.64 14.85 45.74 10.61
GainLoRA (InfLoRA) 53.64 2.87 52.46 4.90
Llama-3-8B O-LoRA wang2023orthogonal 42.49 8.85 38.67 19.28
GainLoRA (O-LoRA) 53.39 3.56 51.69 6.20
InfLoRA liang2024inflora 43.27 6.02 48.77 5.88
GainLoRA (InfLoRA) 52.18 1.40 52.48 4.21
Scaling to Larger Model Architectures To evaluate the effectiveness of our method on larger model architectures, we scale different LoRA-based CL methods to larger models, including T5-XL, Llama-2-7B, Llama-2-13B and Llama-3-8B. Table 2 and Table 3 present the results of different methods. As shown, across models of varying sizes, GainLoRA (O-LoRA) and GainLoRA (InfLoRA) consistently outperform O-LoRA and InfLoRA in terms of AP and FT, respectively. This demonstrates that GainLoRA effectively mitigates forgetting in the new LoRA branch across different model architectures.
Figure 3: (a), (b) and (c) show the number of trainable parameters for different CL methods and model backbones under task sequences Order 1 and Order 2.
Trainable Parameters We compare the number of trainable parameters across different methods for training on different task sequences. The results for T5, Llama-2 and Llama-3 are shown in Figure 3, and the detailed computation of trainable parameters is provided in Appendix B.4. As shown, GainLoRA (O-LoRA) and GainLoRA (InfLoRA) have more trainable parameters than O-LoRA and InfLoRA, respectively. This increase arises from the introduction of the trainable gating module in GainLoRA. However, the additional trainable parameters introduced by GainLoRA are much fewer than those in LoRA. Therefore, the total number of trainable parameters in GainLoRA (O-LoRA) and GainLoRA (InfLoRA) are comparable to that of O-LoRA and InfLoRA, respectively.
Distribution of Outputs in New Gating Module To demonstrate that our GainLoRA effectively minimizes the influence from the new LoRA branches to old tasks, we analyze the output distributions of the new gating modules. Specifically, after training on the final task (i.e., the $15$-th task) in the task sequences, the 15-th task corresponds to the new task, and its associated gating module $g_{15}(\cdot)$ serves as the new gating module.
We obtain the outputs of the new gating module $g_{15}(\cdot)$ on the samples from old and new tasks, respectively. Then, we analyze their distributions in Figure 4. As shown, the outputs of $g_{15}(\cdot)$ for the samples from old tasks are concentrated around 0, effectively minimizing the influence from the new LoRA branch to old tasks. Furthermore, GainLoRA does not constrain the outputs of $g_{15}(\cdot)$ for the samples from the new task. As a result, the outputs of $g_{15}(\cdot)$ for the samples from the new task are distributed near 1, enabling the model to effectively learn the new task.
Figure 4: (a) and (b) show outputs of new gating module in our GainLoRA on different task sequences with T5-Large. (c) and (d) show outputs of new gating module in our GainLoRA on different task sequences with Llama-2-7B.
Table 4: Ablation study of GainLoRA with T5-Large and Llama-2-7B.
Method T5-Large Llama-2-7B
Order 1 Order 2 Order 1 Order 2
[…]
GainLoRA (O-LoRA) 47.84 2.26 46.84 2.91 51.10 4.96 51.14 5.57
No Initialization Constraints 35.30 17.19 39.82 12.90 44.02 11.71 42.89 14.77
No Updating Constraints 23.01 30.32 24.96 28.14 33.74 23.06 34.71 22.36
No Constraints 26.32 26.00 30.63 22.37 34.48 23.46 36.87 21.24
GainLoRA (InfLoRA) 46.21 2.40 46.44 2.61 51.27 2.84 50.17 4.71
No Initialization Constraints 45.38 3.40 43.05 5.15 50.48 3.48 48.17 6.45
No Updating Constraints 37.69 10.94 38.85 9.31 48.52 5.68 47.85 7.00
No Constraints 36.75 12.18 41.00 6.66 49.10 6.07 45.77 8.70
Ablation Study To verify the necessity of both the initialization and updating constraints introduced in Section 3.2.1, we define several variants of GainLoRA. The first variant, referred to as “No Initialization Constraints”, removes the initialization constraints defined in (5). Specifically, it replaces $f(\cdot)$ defined in (8) with function ${\rm sigmoid}(\cdot)$ and eliminates the operation in (10) while keeping all other components unchanged. The second variant, referred to as “No Updating Constraints”, removes the updating constraints defined in (7) by eliminating the operations in (12) while preserving all other components of GainLoRA. The third variant, referred to as “No Constraints”, follows “No Initialization Constraints” and “No Updating Constraints” to remove both the initialization and updating constraints. Table 4 presents the experimental results of these variants. As shown, none of these variants perform as well as our GainLoRA, indicating the critical role of both the initialization constraints and updating constraints in our GainLoRA.
5 Conclusion
In this work, we propose a new method, called GainLoRA, for CL of language models. GainLoRA expands a new LoRA branch for each new task and introduces gating modules to integrate the new and old LoRA branches. Furthermore, GainLoRA leverages the new gating module to minimize the influence of the new LoRA branch to old tasks, effectively mitigating forgetting and improving the model’s overall performance. Experimental results on CL benchmarks demonstrate that GainLoRA outperforms existing state-of-the-art methods.
Limitations Similar to many CL methods for LLMs feng2024tasl ; wang2023orthogonal , our method imposes some constraints on the model to mitigate forgetting. While effective, these constraints may accumulate with increasing tasks, potentially hindering the learning of new tasks. Furthermore, consistent with existing works kirkpatrick2017overcoming ; liang2024inflora ; wang2023orthogonal , our method primarily targets at catastrophic forgetting with non-overlapping tasks, and further investigation is needed to assess its effect on more complex scenarios, such as scenarios where there is overlap between tasks moon2023online ; bang2022online .
6 Broader Impacts
Continual learning (CL) offers a promising direction for improving the efficiency and scalability of language models, particularly in settings with continuously arriving tasks. By enabling incremental updates without retraining from scratch, it significantly reduces computational overhead and resource demands. However, CL often introduces additional components (e.g. memory or gating mechanisms), increasing complexity and requiring effort for maintenance or deployment.
Acknowledgment
This work is supported by NSFC (No.62192783).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims presented in the abstract and introduction accurately reflect the paper’s core contributions and scope. Specifically, the paper clearly positions itself as an improvement over existing LoRA-based methods for CL of language models. The abstract and introduction appropriately emphasize the targeted application domain, namely natural language processing. This alignment ensures that the reader is given a faithful overview of the paper’s goals, proposed methods, and experimental focus.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 5 discusses the limitations of the work performed by the authors.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Yes, the paper provides a complete and correct proof for the theoretical result stated in Proposition 3.1. All necessary assumptions are clearly stated, and the full proof is presented in Section A.3 of the Appendix in the supplemental material. The logical steps are detailed and consistent with the claim, ensuring transparency and rigor in the theoretical justification.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 3 provides a complete description of our method, and Algorithm 1 clearly outlines its step-by-step implementation. Furthermore, Section 4.1 in the main text and Section B of the Appendix present all essential experimental settings, including datasets, evaluation protocols, hyperparameters.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We are committed to ensuring reproducibility and transparency. We have released the full source code along with detailed instructions for reproducing all main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 4.1 in the main text outlines the overall experimental setup, including datasets, evaluation metrics, baseline methods, and training protocols. In addition, Section B of the Appendix provides further details such as data splits, hyperparameter settings (e.g., learning rates, batch sizes), the choice and type of optimizer, and the rationale behind key parameter selections. Together, these sections ensure that the experiments can be accurately reproduced and fairly compared.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: While the main text follows the reporting conventions of existing works on LLMs [75, 15, 61], presenting averaged performance across methods and datasets, we provide more detailed statistical analysis in the Appendix. Specifically, Section C of the supplemental material includes properly defined error bars (e.g., standard deviation across multiple runs), offering insight into the variability and robustness of the reported results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Section B of the Appendix specifies the types of compute resources used, including CPU and GPUs. Additionally, Section B.4 reports the extra computational overhead (in FLOPs) introduced by our method during training and inference.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Section 6 discusses both positive societal impacts and negative societal impacts of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external assets used in this paper, such as T5, Llama-2 and DeepSpeed, are properly credited. We ensure that all assets are used in accordance with their respective licenses and terms of use. Specific citations are provided in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification:
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification:
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A More Details of Methods
A.1 Gradient Projection Memory
We initialize the first $L$ layers of $g_{t}(\cdot)$ using the corresponding layers from the previous gating module $g_{t-1}(\cdot)$. Therefore, the first $L$ layers of $g_{t}(\cdot)$ can be viewed as being initialized at the beginning of the first task and continue their training until the $t$-th task. Additionally, the first $L$ layers in $g_{i}(\cdot)$ serve as checkpoints, preserving the state of $g_{t}(\cdot)$ after learning the $i$-th task ($1\leq i\leq t$). At this time, existing method gradient projection memory (GPM) [47] can be used to learn matrices $\{\bm{M}_{t,l}\}_{l=1}^{L+1}$, where the columns of $\bm{M}_{t,l}$ approximate the orthonormal bases of the subspace $\mathcal{M}_{t,l}$. Specifically, when $t=1$, since there is no old task, ${\mathcal{M}}_{1,l}$ is a null space and $\bm{M}_{1,l}$ is a zero matrix. After learning the $t$-th new task, GPM expands ${\mathcal{M}}_{t,l}$ to ${\mathcal{M}}_{t+1,l}$ by first computing the input matrix $\bm{H}_{t,l}$ where each column of $\bm{H}_{t,l}$ represents an input to the $l$-th layer. Then, the component of $\bm{H}_{t,l}$ already in $\mathcal{M}_{t,l}$ is removed by
$\widehat{\bm{H}}_{t,l}=\bm{H}_{t,l}-\bm{M}_{t,l}(\bm{M}_{t,l})^{T}\bm{H}_{t,l}.$ (15)
Next, singular value decomposition (SVD) is performed on $\widehat{\bm{H}}_{t,l}\widehat{\bm{H}}_{t,l}^{T}$, which is decomposed as $\widehat{\bm{U}}_{t,l}\widehat{\bm{\Sigma}}_{t,l}\widehat{\bm{U}}_{t,l}^{T}$. Then, $u$ new orthonormal bases $\bm{u}_{1},...,\bm{u}_{u}$ are chosen from the columns of $\widehat{\bm{U}}_{t,l}$, where $u$ is the minimum number satisfying the following criteria for a given threshold $\epsilon_{th}$:
[…]
Here, $(\widehat{\bm{H}}_{t,l})_{u}$ denotes the components of $\widehat{\bm{H}}_{t,l}$ corresponding to the top-$u$ singular values. Then, the orthonormal bases of subspace ${\mathcal{M}}_{t+1,l}$ are obtained by augmenting the orthonormal bases of subspace $\mathcal{M}_{t,l}$ with the new orthogonal vectors $\bm{u}_{1},...,\bm{u}_{u}$, resulting in $\bm{M}_{t+1,l}=[\bm{M}_{t,l},\bm{u}_{1},...,\bm{u}_{u}]$.
A.2 More Details of O-LoRA and InfLoRA
O-LoRA
O-LoRA [61] ensures that the new LoRA branch remains orthogonal to all the old LoRA branches. Specifically, during the learning of the $t$-th new task with the $t$-th LoRA branch ($\bm{A}_{t},\bm{B}_{t}$), O-LoRA computes the inner product between the new and old LoRA branches as
$\bm{O}_{i,t}=\bm{B}_{i}^{T}\bm{B}_{t}~~~{\rm for}~1\leq i\leq t-1$ (17)
Then, the loss function of O-LoRA is defined as
[…]
For further details on O-LoRA, we refer readers to the original paper [61].
InfLoRA
InfLoRA [32] ensures orthogonality between the new LoRA branch and the gradients of old tasks. Specifically, it shows that only fine-tuning the down-projection matrix $\bm{A}_{t}$ in the new LoRA branch is equivalent to directly fine-tuning the pre-trained weights within a subspace spanned by the rows of $\bm{B}_{t}$. Therefore, before learning the $t$-th task, InfLoRA designs $\bm{B}_{t}$ to be orthogonal to the gradients of the old tasks. During the learning of the $t$-th task, InfLoRA only tunes $\bm{A}_{t}$ in the new LoRA branch while freezing $\bm{B}_{t}$ and all the old LoRA branches. For further details on InfLoRA, we refer readers to the original paper [32].
A.3 Proof of Proposition 3.1
Proposition A.1.
If the constraints in (7) are satisfied, subspaces $\{\mathcal{M}_{t,l}\}_{l=1}^{L+1}$ remain unchanged during the learning of the $t$-th task. Furthermore, for any input $\bm{x}$ from the previous t1t-1 tasks, $g_{t}(\bm{x})$ remains unchanged during the learning of the $t$-th task.
Proof.
For any $\bm{x}$ from previous t1t-1 tasks, we rewrite $\bm{g}_{t}(\bm{x})$ as
$g_{t}(\bm{x})=f(\bm{G}_{t,L+1}\bm{p}_{t,L}),$
$\bm{p}_{t,l}=\sigma(\bm{G}_{t,l}\bm{p}_{t,l-1}),~l\in\{1,2,...,L\},$
$\bm{p}_{t,0}=\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})).$ (19)
Since $\bm{p}_{t,0}={\rm Pool}({\rm Token}(\bm{x}))$ is unrelated to the parameters of the new gating module $g_{t}(\cdot)$, $\bm{p}_{t,0}$ does not change with the update of $g_{t}(\cdot)$. Since $\mathcal{M}_{t,1}$ is spanned by $\bm{p}_{t,0}$, $\mathcal{M}_{t,1}$ remains unchanged during the learning of the $t$-th task.
Suppose that we have proven that $\bm{p}_{t,l-1}$ does not change with the update of the new gating module $g_{t}(\cdot)$ ($1\leq l\leq L$). Since $\mathcal{M}_{t,l}$ is spanned by $\bm{p}_{t,l-1}$, $\mathcal{M}_{t,l}$ remains unchanged during the learning of the $t$-th task. At this point, $\bm{p}_{t,l}$ can be expressed as
$\bm{p}_{t,l}=\sigma(({\rm Init}(\bm{G}_{t,l})+\Delta\bm{G}_{t,l})\bm{p}_{t,l-1})=\sigma({\rm Init}(\bm{G}_{t,l})\bm{p}_{t,l-1}).$ (20)
Here, the second equality holds since $\bm{p}_{t,l-1}\in\mathcal{M}_{t,l}$ and $\Delta\bm{G}_{t,l}\bot\mathcal{M}_{t,l}$. Therefore, $\bm{p}_{t,l}$ does not change with the update of the new gating module $g_{t}(\cdot)$ ($1\leq l\leq L$). Since $\mathcal{M}_{t,l+1}$ is spanned by $\bm{p}_{t,l}$, $\mathcal{M}_{t,l+1}$ remains unchanged during the learning of the $t$-th task.
Furthermore, during the learning of the $t$-th task, $g_{t}(\bm{x})$ can be expressed as
$\bm{g}_{t}(\bm{x})=f(({\rm Init}(\bm{G}_{t,L+1})+\Delta\bm{G}_{t,L+1})\bm{p}_{t,L})=f({\rm Init}(\bm{G}_{t,L+1})\bm{p}_{t,L}).$ (21)
Here, the second equality holds since $\bm{p}_{t,L}\in\mathcal{M}_{t,L+1}$ and $\Delta\bm{G}_{t,L+1}\bot\mathcal{M}_{t,L+1}$. ∎
Appendix B More Details of Experimental Settings
B.1 More Details of Datasets
Table 5 and Table 6 show the details of Long Sequence Benchmark and SuperNI Benchmark, respectively. Long Sequence Benchmark consists of 15 classification tasks while SuperNI Benchmark consists of various NLP tasks, including dialogue generation, information extraction, question answering, summarization, and sentiment analysis.
Table 5: Details of different tasks in Long Benchmark.
Dataset name Category Domain Task Type Metric
Yelp CL Benchmark sentiment analysis Yelp reviews Accuracy
Amazon CL Benchmark sentiment analysis Amazon reviews Accuracy
DBpedia CL Benchmark topic classification Wikipedia Accuracy
Yahoo CL Benchmark topic classification Yahoo Q&A Accuracy
AG News CL Benchmark topic classification news Accuracy
MNLI GLUE natural language inference various Accuracy
QQP GLUE paraphrase detection Quora Accuracy
RTE GLUE natural language inference news, Wikipedia Accuracy
SST-2 GLUE sentiment analysis movie reviews Accuracy
WiC SuperGLUE word sense disambiguation lexical databases Accuracy
CB SuperGLUE natural language inference various Accuracy
COPA SuperGLUE question and answering blogs, encyclopedia Accuracy
BoolQA SuperGLUE boolean question and answering Wikipedia Accuracy
MultiRC SuperGLUE question and answering various Accuracy
IMDB SuperGLUE sentiment analysis movie reviews Accuracy
Table 6: Details of different tasks in SuperNI Benchmark.
Dataset name Task Type Metric
Task639_\_multi_\_woz_\_user_\_utterance_\_generation summarization Rouge-L
Task1590_\_diplomacy_\_text_\_generation summarization Rouge-L
Task1729_\_personachat_\_generate_\_next summarization Rouge-L
Task181_\_outcome_\_extraction information extraction Rouge-L
Task748_\_glucose_\_reverse_\_cause_\_event_\_detection information extraction Rouge-L
Task1510_\_evalution_\_relation_\_extraction information extraction Rouge-L
Task002_\_quoref_\_answer_\_generation dialogue generation Rouge-L
Task073_\_commonsenseqa_\_answer_\_generation dialogue generation Rouge-L
Task591_\_sciq_\_answer_\_generation dialogue generation Rouge-L
Task511_\_reddit_\_tifu_\_long_\_text_\_summarization question answering Rouge-L
Task1290_\_xsum_\_summarization question answering Rouge-L
Task1572_\_samsum_\_summary question answering Rouge-L
Task363_\_sst2_\_polarity_\_classification sentiment analysis Accuracy
Task875_\_emotion_\_classification sentiment analysis Accuracy
Task1687_\_sentiment140_\_classification sentiment analysis Accuracy
The task sequences are constructed using Long Sequence Benchmark and SuperNI Benchmark. The details of different task sequences are presented in Table 7.
Table 7: The order of different task sequences for experiments.
Benchmark Order Task Sequence
SuperNI Benchmark 1
task1572 $\rightarrow$ task363 $\rightarrow$ task1290 $\rightarrow$ task181 $\rightarrow$ task002 $\rightarrow$
task1510 $\rightarrow$ task639 $\rightarrow$ task1729 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$
task748 $\rightarrow$ task511 $\rightarrow$ task591 $\rightarrow$ task1687 $\rightarrow$ task875
2
task748 $\rightarrow$ task073 $\rightarrow$ task1590 $\rightarrow$ task639 $\rightarrow$ task1572 $\rightarrow$
task1687 $\rightarrow$ task591 $\rightarrow$ task363 $\rightarrow$ task1510 $\rightarrow$ task1729 $\rightarrow$
task181 $\rightarrow$ task511 $\rightarrow$ task002 $\rightarrow$ task1290 $\rightarrow$ task875
CL Benchmark 3
MNLI $\rightarrow$ CB $\rightarrow$ WiC $\rightarrow$ COPA $\rightarrow$ QQP $\rightarrow$
BoolQA $\rightarrow$ RTE $\rightarrow$ IMDB $\rightarrow$ Yelp $\rightarrow$ Amazon $\rightarrow$
SST-2 $\rightarrow$ DBpedia $\rightarrow$ AG News $\rightarrow$ MultiRC $\rightarrow$ Yahoo
4
Yelp $\rightarrow$ Amazon $\rightarrow$ MNLI $\rightarrow$ CB $\rightarrow$ COPA $\rightarrow$
QQP $\rightarrow$ RTE $\rightarrow$ IMDB $\rightarrow$ SST-2 $\rightarrow$ DBpedia $\rightarrow$
AG News $\rightarrow$ Yahoo $\rightarrow$ MultiRC $\rightarrow$ BoolQA $\rightarrow$ WiC
B.2 More Implementation Details
Following existing CL works [40, 61, 68], all methods are implemented using instruction tuning [40]. Experiments are conducted on 5 NVIDIA RTX A6000 GPUs with AdamW [35] as the optimizer. The type of CPU is Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz. For T5-Large and T5-XL, their relatively smaller model sizes allow experiments to be performed on a single A6000 GPU with gradient accumulation. For Llama-2-7B and Llama-2-13B, data parallelism with DeepSpeed ZeRO-2 [44] is prioritized across multiple A6000 GPUs. FlashAttention-2 [9] is employed to reduce memory usage during training, ensuring sufficient GPU memory to enable DeepSpeed ZeRO-2 whenever possible. However, if the sequence lengths of certain tasks are too long to enable DeepSpeed ZeRO-2 even with FlashAttention-2, DeepSpeed ZeRO-3 is utilized to handle these tasks.
To ensure fair comparisons, for all the methods based on LoRA, we follow existing CL methods [21, 61, 75] by integrating the LoRA architecture into the query and value components of the multi-head attention mechanism in each Transformer block. Following existing works [61, 75], for all the methods based on LoRA, the rank of a single LoRA branch is set to 4 for Order 1 and Order 2, and 8 for Order 3 and Order 4. We also vary the rank in LoRA branches and show the results in Appendix C.4.
For our methods, the global batch size is set to 32 across all model backbones. The learning rate is set to 3e-4 for T5 backbones and 5e-5 for Llama backbones. Each task is trained for 100 epochs with T5 backbones and 50 epochs with Llama backbones. For baselines, we follow their official implementations to set the hyperparameters, making the comparison as fair as possible. If this does not achieve the expected performance, we perform a hyperparameter search for the learning rate and batch size.
B.3 More Details about the Architecture of the Gating Module
The architecture of the gating module $g_{i}(\cdot)$ can be represented as
$g_{i}(\bm{x})=f(\bm{G}_{i,L+1}\bm{p}_{i,L}),$
$\bm{p}_{i,l}=\sigma(\bm{G}_{i,l}\bm{p}_{i,l-1}),~l\in\{1,2,...,L\},$
$\bm{p}_{i,0}=\bm{p}_{0}={\rm Pool}({\rm Token}(\bm{x})).$ (22)
Non-linear activation function $\sigma(\cdot)$ is set to SiLU [13]. For all experiments, unless otherwise stated, $L$ is set to 2. In other words, the gating module $g_{i}(\cdot)$ has three layers. For T5-Large and T5-XL, the parameters in the $i$-th gating module $g_{i}(\cdot)$ are $\bm{G}_{i,1}\in\mathbb{R}^{100\times d}$, $\bm{G}_{i,2}\in\mathbb{R}^{d\times 100}$ and $\bm{G}_{i,3}\in\mathbb{R}^{1\times d}$. For Llama-2-7B and Llama-2-13B, the parameters in the $i$-th gating module $g_{i}(\cdot)$ are $\bm{G}_{i,1}\in\mathbb{R}^{50\times d}$, $\bm{G}_{i,2}\in\mathbb{R}^{d\times 50}$ and $\bm{G}_{i,3}\in\mathbb{R}^{1\times d}$. Here, $d$ denotes the dimension of the embeddings. For different models, $d$ is 1024 for T5-Large and T5-XL, 4096 for Llama-2-7B, and 5120 for Llama-2-13B.
Additionally, we investigate the influence of the architecture of the gating module on the performance of our method. Results are provided in Appendix C.3.
B.4 Computation of Trainable Parameters
To ensure fair comparisons, we set the same rank for each LoRA branch across all CL methods based on the expandable LoRA architectures shown in Figure 1 (a). Additionally, for all the methods based on LoRA, the LoRA modules are incorporated into the query and value components of the multi-head attention mechanism within each Transformer block.
B.4.1 Computation of Trainable Parameters in T5-Large
In T5-Large, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{1024\times 1024}$. The model consists of 24 self-attention modules in the encoder, 24 self-attention modules in the decoder, and 24 cross-attention modules in the decoder, resulting in a total of (24+24+24)2=144(24+24+24)*2=144 pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 1024}$, resulting in 1024r144+r1024144=294912r1024*r*144+r*1024*144=294912r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 2949124=1179648=1.18294912*4=1179648=1.18M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$, resulting in 1024r144=147456r1024*r*144=147456r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 147456r=589824=0.59147456r=589824=0.59M.
GainLoRA introduces an additional new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{100\times 1024}$, $\bm{G}_{t,2}\in\mathbb{R}^{1024\times 100}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 1024}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 1179648+1024100+1024100+1024=1385472=1.391179648+1024*100+1024*100+1024=1385472=1.39M. The number of trainable parameters in GainLoRA (InfLoRA) is 589824+1024100+1024100+1024=795648=0.80589824+1024*100+1024*100+1024=795648=0.80M.
B.4.2 Computation of Trainable Parameters in T5-XL
In T5-XL, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{4096\times 1024}$. The model architecture is similar to T5-Large, with 144 pre-trained weights incorporating LoRA.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 1024}$, resulting in is 4096r144+r1024144=737280r4096*r*144+r*1024*144=737280r trainable parameters. When $r=4$, O-LoRA has 7372804=2949120=2.95737280*4=2949120=2.95M trainable parameters. InfLoRA only updates $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$, resulting in 4096r144=589824r4096*r*144=589824r trainable parameters. When $r=4$, InfLoRA has 5898244=2359296=2.36589824*4=2359296=2.36M trainable parameters.
GainLoRA introduces the same new gating module $g_{t}(\cdot)$ as in T5-Large, with parameters $\bm{G}_{t,1}\in\mathbb{R}^{100\times 1024}$, $\bm{G}_{t,2}\in\mathbb{R}^{1024\times 100}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 1024}$. Thus, the total number of trainable parameters for GainLoRA (O-LoRA) is 2949120+1024100+1024100+1024=3154944=3.152949120+1024*100+1024*100+1024=3154944=3.15M. The total number of trainable parameters in GainLoRA (InfLoRA) is 2359296+1024100+1024100+1024=2565120=2.572359296+1024*100+1024*100+1024=2565120=2.57M.
B.4.3 Computation of Trainable Parameters in Llama-2-7B
In Llama-2-7B, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{4096\times 4096}$. The model contains 32 self-attention modules, resulting in 322=6432*2=64 pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$, resulting in 4096r64+r409664=524288r4096*r*64+r*4096*64=524288r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 5242884=2097152=2.10524288*4=2097152=2.10M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$, resulting in 4096r64=262144r4096*r*64=262144r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 2621444=1048576=1.05262144*4=1048576=1.05M.
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 4096}$, $\bm{G}_{t,2}\in\mathbb{R}^{4096\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 4096}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 2097152+409650+409650+4096=2510848=2.512097152+4096*50+4096*50+4096=2510848=2.51M. The number of trainable parameters in GainLoRA (InfLoRA) is 1048576+409650+409650+4096=1462272=1.461048576+4096*50+4096*50+4096=1462272=1.46M.
B.4.4 Computation of Trainable Parameters in Llama-2-13B
In Llama-2-13B, the projection weights for the query and value components have shapes $\bm{W}_{q},\bm{W}_{v}\in\mathbb{R}^{5120\times 5120}$. The model contains 40 self-attention modules, resulting in 402=8040*2=80 pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{5120\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 5120}$, resulting in 5120r80+r512080=819200r5120*r*80+r*5120*80=819200r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 8192004=3276800=3.28819200*4=3276800=3.28M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{5120\times r}$, resulting in 5120r80=409600r5120*r*80=409600r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 4096004=1638400=1.64409600*4=1638400=1.64M.
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 5120}$, $\bm{G}_{t,2}\in\mathbb{R}^{5120\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 5120}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 3276800+512050+512050+5120=3793920=3.793276800+5120*50+5120*50+5120=3793920=3.79M. The number of trainable parameters in GainLoRA (InfLoRA) is 1638400+512050+512050+5120=2155520=2.161638400+5120*50+5120*50+5120=2155520=2.16M.
B.4.5 Computation of Trainable Parameters in Llama-3-8B
In Llama-3-8B, the projection weights for the query and value components have shapes $\bm{W}_{q}\in\mathbb{R}^{4096\times 4096},\bm{W}_{v}\in\mathbb{R}^{4096\times 1024}$. The model contains 40 self-attention modules, resulting in 322=6432*2=64 pre-trained weights that incorporate the LoRA architecture.
During the learning of the $t$-th new task, O-LoRA updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$ for query and $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ and $\bm{B}_{t}\in\mathbb{R}^{r\times 4096}$ for value, resulting in 1024r32+4096r32+r409664=425984r1024*r*32+4096*r*32+r*4096*64=425984r trainable parameters. When $r=4$, the number of trainable parameters in O-LoRA is 4259844=1703936=1.70425984*4=1703936=1.70M. InfLoRA only updates the parameters $\bm{A}_{t}\in\mathbb{R}^{4096\times r}$ for query and $\bm{A}_{t}\in\mathbb{R}^{1024\times r}$ for value, resulting in 4096r32+1024r32=163840r4096*r*32+1024*r*32=163840r trainable parameters. When $r=4$, the number of trainable parameters in InfLoRA is 1638404=655360=0.66163840*4=655360=0.66M.
GainLoRA introduces a new gating module $g_{t}(\cdot)$ with parameters $\bm{G}_{t,1}\in\mathbb{R}^{50\times 4096}$, $\bm{G}_{t,2}\in\mathbb{R}^{4096\times 50}$ and $\bm{G}_{t,3}\in\mathbb{R}^{1\times 4096}$. Therefore, the number of trainable parameters in GainLoRA (O-LoRA) is 1703936+409650+409650+4096=2117632=2.121703936+4096*50+4096*50+4096=2117632=2.12M. The number of trainable parameters in GainLoRA (InfLoRA) is 655360+409650+409650+4096=1069056=1.07655360+4096*50+4096*50+4096=1069056=1.07M.
Table 8: FLOPs and MACs for different models.
Method Input Shape (batch,length) FLOPs (G) MACs (G)
T5-Large Original (1,128) 194.25 97.1
GainLoRA (O-LoRA) (1,128) 198.79 99.37
GainLoRA (InfLoRA) (1,128) 198.79 99.37
T5-XL Original (1,128) 751.7 375.78
GainLoRA (O-LoRA) (1,128) 763.03 381.45
GainLoRA (InfLoRA) (1,128) 763.03 381.45
Llama-2-7B Original (1,128) 1701.07 850.5
GainLoRA (O-LoRA) (1,128) 1709.14 854.53
GainLoRA (InfLoRA) (1,128) 1709.14 854.53
Llama-2-13B Original (1,128) 3291.66 1645.79
GainLoRA (O-LoRA) (1,128) 3304.26 1652.09
GainLoRA (InfLoRA) (1,128) 3304.26 1652.09
Llama-3-8B Original (1,128) 1929.86 964.89
GainLoRA (O-LoRA) (1,128) 1930.49 965.21
GainLoRA (InfLoRA) (1,128) 1930.49 965.21
Appendix C More Experimental Results
C.1 Discussing Computational Costs Introduced by GainLoRA
Existing methods, such as O-LoRA and InfLoRA, adopt the expandable LoRA architecture shown in Figure 1 (a) and fix the integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1, allowing the model to merge the expanded LoRA branches into the pre-trained matrix at inference time, thereby avoiding additional computational costs. However, when using our gating module to integrate different LoRA branches, the LoRA branches cannot be merged into the pre-trained matrix at inference time, which introduces additional computational costs. Nevertheless, we demonstrate that these computational costs are minimal compared to the computational cost of the original large language models (LLMs).
Table 8 compares the floating-point operations (FLOPs) and multiply-add operations (MACs) during inference for different models with and without GainLoRA. The computation of FLOPs and MACs follows the existing project calflops [69]. Here, “Original” denotes the original LLMs without any LoRA adaptation. Methods such as O-LoRA and InfLoRA avoid additional computational costs by merging their LoRA branches into the original weights during inference, resulting in FLOPs and MACs identical to the original LLMs. Despite introducing additional FLOPs and MACs compared to the original LLMs, GainLoRA maintains minimal computational overhead relative to the original LLMs.
C.2 Additional Computation Introduced by Subspace Construction
The memory and computational overhead of subspace construction in GainLoRA is minimal due to the small size of the gating module (only 3 layers, see B.3). We provide detailed analyses below.
[…]
Computational Overhead The computational overhead for subspace construction requires a single forward pass over the task dataset and SVD on the feature matrices of the gating module.
[…]
GPM requires performing SVD on matrix $H_{l}H_{l}^{T}\in\mathbb{R}^{d_{l}\times d_{l}}$, where $H_{l}$ is the feature matrix in the $l$-th layer of gating module. Based on existing conclusions [55], the FLOPs for SVD on $HH^{T}$ is less than $10d_{l}^{3}$. For T5-Large ($d_{1}=d_{3}=1024$ and $d_{2}=100$), this results in $10*(2*1024^{3}+100^{3})<30GFLOPs$, which is negligible compared to a single forward pass with sequence length 128 (see Table 8). Similar calculations give the same conclusion for T5-XL, Llama-2-7B, Llama-2-13B and Llama-3-8B.
C.3 Varying the Architecture of Gating Module
C.3.1 Varying Function $f(\cdot)$ in Gating Module
To implement our method, we define function $f(\cdot)$ as (8). Here, we vary the formula of function $f(\cdot)$ as the following two functions:
${\rm min}\{|b|,1\},~|\sin(\frac{\pi b}{2})|.$ (23)
Clearly, these two functions map real values among $[0,1]$ and satisfy $f(0)=0$. Table 9 shows the results. As we can see, when changing the formula of $f(\cdot)$, GainLoRA also improves the performance of O-LoRA and InfLoRA.
Table 9: Varying the function $f(\cdot)$ in GainLoRA on different task sequences with T5-large model.
Method Order 1 Order 2
[…]
GainLoRA (InfLoRA) $\left(f(b)=|2{\rm sigmoid}(b)-1|\right)$ 46.21 2.40 46.44 2.61
GainLoRA (InfLoRA) $\left(f(b)={\rm min}\{|b|,1\}\right)$ 45.05 2.07 45.00 1.74
GainLoRA (InfLoRA) $\left(f(b)=|\sin(\frac{\pi b}{2})|\right)$ 47.48 1.21 45.03 2.37
InfLoRA 39.78 7.64 39.57 8.93
GainLoRA (O-LoRA) $\left(f(b)=|2{\rm sigmoid}(b)-1|\right)$ 47.84 2.26 46.84 2.91
GainLoRA (O-LoRA) $\left(f(b)={\rm min}\{|b|,1\}\right)$ 49.62 2.83 48.62 3.74
GainLoRA (O-LoRA) $\left(f(b)=|\sin(\frac{\pi b}{2})|\right)$ 48.49 3.84 47.20 4.69
O-LoRA 26.37 19.15 32.83 11.99
C.3.2 Varying the Shapes of Weights in Gating Module
In this section, we vary the shapes of the weights in the gating modules with T5-Large. Specifically, we set the weights $\bm{G}_{i,1}\in\mathbb{R}^{d_{h}\times 1024}$ and $\bm{G}_{i,2}\in\mathbb{R}^{1024\times d_{h}}$ in each gating module $g_{i}(\cdot)$ and vary $d_{h}$ over $\{50,100,200\}$. Figure 5 (a) and Figure 5 (b) show the results. As we can see, when increasing $d_{h}$, the performance of GainLoRA remains relatively stable, indicating that our method is robust to the shape of the weights in the gating module. Note that the number of trainable parameters increases as $d_{h}$ increases.
Figure 5: (a) and (b) show the variation of our methods’ performance with the shapes of the weights in the gating module. (c) and (d) show the variation of our methods’ performance with the Layers of the gating module.
C.3.3 Varying the Layers of Gating Module
In this section, we vary the layers of the gating modules with T5-Large. Specifically, we vary across $\{0,2,4\}$. when $L=0$, there is only one layer with $\bm{G}_{i,1}\in\mathcal{R}^{1\times 1024}$ in each gating module $g_{i}(\cdot)$. When $L=2$, there are three layers with $\bm{G}_{i,1}\in\mathcal{R}^{100\times 1024}$, $\bm{G}_{i,2}\in\mathcal{R}^{1024\times 100}$ and $\bm{G}_{i,3}\in\mathcal{R}^{1\times 1024}$. When $L=4$, there are 5 layers with $\bm{G}_{i,1}\in\mathcal{R}^{100\times 1024}$, $\bm{G}_{i,2}\in\mathcal{R}^{1024\times 100}$, $\bm{G}_{i,3}\in\mathcal{R}^{100\times 1024}$, $\bm{G}_{i,4}\in\mathcal{R}^{1024\times 100}$, and $\bm{G}_{i,5}\in\mathcal{R}^{1\times 1024}$ in each gating module. Figure 5 (c) and Figure 5 (d) show the results. As we can see, when increasing the layers of gating modules, the performance of GainLoRA remains relatively stable, indicating that our method is robust to the layers of the gating module. Note that the number of trainable parameters increases as the number of layers in gating modules increases.
Figure 6: The variation of our methods’ performance with the Layers of the gating module.
C.4 Varying Ranks in LoRA Branches
In this section, we vary the rank of LoRA branches across $\{2,4,8\}$ with T5-Large. Figure 6 shows the results. As shown, when the rank of LoRA branches increases, the performance of GainLoRA remains relatively stable. Note that the number of trainable parameters increases as the rank of LoRA branches increases.
C.5 Adopting Other Update Strategies for the New LoRA Branch
Our GainLoRA does not impose specific update strategies for the new LoRA branches. In this work, we adopt the same update strategies as the existing two methods, O-LoRA [61] and InfLoRA [32]. Related methods, such as IncLoRA [21] and C-LoRA [50], also adopt the expandable LoRA architecture illustrated in Figure 1 (a) and fix all integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1. Our method GainLoRA can also adopt their update strategies for the new LoRA branch. Table 10 presents the results, demonstrating that GainLoRA further improves the performance of these two methods.
Table 10: The overall results on different task sequences with T5-large model.
Method Order 1 Order 2
[…]
IncLoRA 12.33 41.93 16.65 36.56
GainLoRA (IncLoRA) 47.82 3.73 45.42 5.83
C-LoRA 22.69 24.25 32.81 11.60
GainLoRA (C-LoRA) 49.24 2.94 46.23 6.05
C.6 Performance on the TRACE Benchmark
To further demonstrate our method’s effectiveness, we follow existing CL methods for LLMs [18, 62] and conduct experiments on the TRACE dataset [62] with Llama-2-7B-Chat. The dataset comprises a diverse set of challenging instruction-tuned tasks, spanning multilingual comprehension, domain-specific knowledge, arithmetic reasoning, and coding. An overview of the tasks in TRACE is provided in Table 11.
Table 11: The order of TRACE benchmark for experiments.
Benchmark Task Sequence
TRACE Benchmark
C-STANCE $\rightarrow$ FOMC $\rightarrow$ MeetingBank $\rightarrow$ Py150 $\rightarrow$
ScienceQA $\rightarrow$ NumGLUE-cm $\rightarrow$ NUMGLUE-ds $\rightarrow$ 20Minuten
Table 12: TRACE benchmark performance using LLama-2-7B-Chat.
[…]
O-LoRA [61] 41.04 8.05
GainLoRA (O-LoRA) 48.10 0.99
InfLoRA [32] 47.67 2.25
GainLoRA (InfLoRA) 49.15 0.89
Table 13: Comparison of general ability scores across six diverse evaluation tasks between the base LLaMA-2-7B chat model and different methods.
PIQA MMLU GSM8K BBH BoolQA TydiQA
O-LoRA [61] 72.85 32.87 13.42 35.10 56.88 19.48
GainLoRA (O-LoRA) 73.61 33.33 18.57 36.47 59.69 25.00
InfLoRA [32] 74.86 40.86 15.69 35.87 65.29 27.25
GainLoRA (InfLoRA) 75.24 44.25 21.30 37.44 68.81 27.84
Llama-2-7B-Chat 75.35 46.13 26.54 40.09 70.46 23.45
Table 12 reports the average performance on the TRACE benchmark after sequentially learning all tasks. The results demonstrate that our method effectively mitigates catastrophic forgetting and outperforms existing baselines. This capability is crucial for real-world applications.
Retention of General Capabilities
We also follow existing work [62] to explicitly evaluate the preservation of general abilities, such as instruction-following, after continual learning on the TRACE benchmark using different methods. Table 13 indicates that continual learning with different methods often leads to a degradation of general abilities. However, GainLoRA demonstrates a stronger ability to mitigate forgetting compared to other LoRA-based methods, including O-LoRA and InfLoRA.
C.7 Compared with More CL Methods in CV
Following many existing continual learning methods in NLP [75, 61, 18], this paper focuses on models based on next-token prediction, such as T5 [43] and LLaMA [54], which lack the [CLS] token used in ViT. Although many continual learning methods based on pre-trained models in computer vision [66, 67, 63, 57, 58, 51] cannot be directly applied to our setting, we adapt several of them to the T5 architecture to ensure a comprehensive comparison. Specifically, We implement these methods by injecting prompts into both the keys and values in T5, and introduce zero-padding in the position bias tensor to ensure shape compatibility. Note that we do not add positional information to the prompts, which is consistent with DualPrompt and CODA-Prompt in ViT. The results on order 1 with T5 are reported in Table 14, showing that these methods perform significantly worse than our GainLoRA and exhibit noticeable forgetting.
Table 14: Compare with different methods on order 1 with T5 architecture.
[…]
L2P [67] 15.23 11.34
DualPrompt [66] 17.40 10.63
CODA-Prompt [51] 19.28 14.62
GainLoRA (O-LoRA) 47.84 2.26
GainLoRA (InfLoRA) 46.21 2.40